find three consecutive integers such that the sum of the first two integers is 24 more than the third integer.
Answer by checkley71(8403)

LET X, X+1 & X+2 BE THE THREE CONSECUTIVE INTEGERS.
X+(X+1)=24+(X+2)
2X+1=24+X+2
2X-X=24+2-1
X=25 ANSWER FOR THE FIRST INTEGER.
25+1=26 FOR THE SECOND
25+2=27 FOR THE THIRD.
PROOF
25+26=24+27
51=51